OpenStudy (lexy148):

How much interst will be earned at the end of five years on a deposit of $2,500 in anaccount paying 4.25% interest compounded monthly?

OpenStudy (anonymous):

\[2500\times (1+\frac{.0425}{12})^{60}\] and a calculator

OpenStudy (lexy148):

i got 3090.75

OpenStudy (lexy148):

idk if thats correct

OpenStudy (lexy148):

@satellite73

OpenStudy (anonymous):

i didn't do it give me a second

OpenStudy (lexy148):

ok

OpenStudy (anonymous):

of course if the question really just asks for the interest, not the total, then you have to subtract \(2500\) from that total

OpenStudy (anonymous):

oh and yeah, your total was right

OpenStudy (lexy148):

3090.75-2500= 590.75

OpenStudy (anonymous):

i believe you

OpenStudy (lexy148):

so that would be my interst

OpenStudy (anonymous):

yes

OpenStudy (lexy148):

ok thank you